OpenStudy (anonymous1):

Show that the graphs of r = aθ and rθ = a have an unlimited number of intersection points.

OpenStudy (anonymous1):

I know that the equation \[r=a\theta\] has the same graph as \[-r=a(\theta+\pi)\] and that the equation \[r\theta=a\] has the same graph as \[-r(\theta+\pi)=a\]